Balance the following redox equations by the ion-electron half-reaction method:

(a) (acid solution): BrO3 – + Cr3+ → Br2 + HCrO4 –

(b) (acid solution): MnO4 – + VO2+ → Mn2+ + V(OH)4 +

(c) (basic solution): MnO4 – + OH– → MnO4 2– + O2 (g)
Answer #1

Balanced equations are

(a) (Acid soluition):  6BrO3 + 10Cr3+ +22H2O =   3Br2 + 10HCrO4- +34H+

(b) (Acid solution):  MnO-4 + 5VO2+ +H2O = Mn2+ + 5VO2+2 +2H+

(c)  (Basic solution):  MnO-4 + 4OH- = MnO2-4 + O2 + 2H2O